13/11y - 2/11y = -3
Answer by checkley77(12844)

13y/11 - 2y/11 = -3
(13y-2y)/11=-3
11y/11=-3
-3*11=11y
11y=-33
y=-33/11
y=-3 ans.
Proof:
13/11*-3-2/11*-3=-3
-39/11+6/11=-3
-33/11=-3
-3=-3